Two bodies of equal masses are in uniform circular motion with the same period.If r1 and r2 are the radii of their circular path the ratio of their centripetal force is ___?

Grade:11

1 Answers

DR STRANGE
7 years ago
HI ARUN,
SO YOU ARE CONFUSED BECAUSE THERE TWO FORMULA FOR CENTRIPETAL FORCE
F= MRW F= MV2/R
IN THIS CASE YOU WILL USE FIRST ONE AS THE MASSES ARE SAME AND THE TIME PERIOD i.e. W (ANGULAR VELOCITY) IS SAME.
SO BY JUST TAKING THE RATIO F1/F2 =MR1W2/MR2W= R1/R2.
HAVE A GOOD DAY :)